OpenStudy (anonymous):

can anyone write me the process to of finding the int. sin^2(x)cos^4(x)dx?

OpenStudy (anonymous):

\[\int\limits \sin^2xcos^4xdx\]

OpenStudy (anonymous):

\[\sin ^2(x) \cos ^4(x)=\frac{1}{32} (\cos (2 x)-2 \cos (4 x)-\cos (6 x)+2) \]

OpenStudy (anonymous):

\[\int\limits \sin ^2(x) \cos ^4(x)dx=\frac{x}{16}+\frac{1}{64} \sin (2 x)-\frac{1}{64} \sin (4 x)-\frac{1}{192} \sin (6 x) \]